### v1

@@ Abstract
Event cameras, or Dynamic Vision Sensors (DVS) are novel neuromorphic sensors that capture brightness changes as a continuous stream of “events” rather than traditional intensity frames. Converting sparse events to dense intensity frames faithfully has long been an ill-posed problem. Previous methods have primarily focused on converting events to video in dynamic scenes or with a moving camera. In this paper, for the first time, we realize events to dense intensity image conversion using a stationary event camera in static scenes. Different from traditional methods that mainly rely on event integration, the proposed Event-Based Temporal Mapping Photography (EvTemMap) measures the time of event emitting for each pixel. Then, the resulting Temporal Matrix is converted to an intensity frame with a temporal mapping neural network. At the hardware level, the proposed EvTemMap is implemented by combining a transmittance adjustment device with a DVS, named Adjustable Transmittance Dynamic Vision Sensor. Additionally, we collected TemMat dataset under various conditions including low-light and high dynamic range scenes. The experimental results showcase the high dynamic range, fine-grained details, and high-grayscale-resolution of the proposed EvTemMap, as well as the enhanced performance on downstream computer vision tasks compared to other methods. The code and TemMat dataset will be made publicly available.
@@
Event-based vision Dynamic Vision Sensor Image reconstruction High dynamic range imagingLow-light photography
@@ 1 Introduction
Event cameras offer high temporal resolution and dynamic range, swiftly responding to brightness changes through positive and negative events. While the distinctive asynchronous data output from event cameras diminishes data redundancy, it concurrently discards the grayscale information of the image. To capture grayscale images, a dynamic and active pixel vision sensor (DAVIS) [5] integrates an active pixel sensor (APS) and a DVS sensor on a single chip. This integration allows the sensor to output both grayscale frames and events. The fusion of grayscale images with events facilitates various computer vision tasks. Many event-based image/video enhancement tasks rely on such devices, such as motion deblur [37], video interpolation [38], high dynamic range (HDR) imaging [27], auto-focusing [3] etc. These tasks leverage the high temporal resolution and dynamic range of events to enhance grayscale images.
@@
The temporal dimension of events exhibits theoretically infinite description levels, with 1 $\mu$s as one level. Can we utilize time to measure grayscale? Imagine a scene: the event camera is initially in complete darkness, and when the lighting gradually increases, the event camera captures positive events as each pixel gets brighter. Each pixel lights up at a different rate, depending on its corresponding point’s light intensity in the scene. As the event sensor initiates from darkness and gradually intensifies, the time each pixel takes to reach a specified threshold (i.e., the timestamp of the initial positive event, IPE) indicates its brightening speed. Hence, through the measurement of the timestamp associated with each IPE, we can derive the light intensity of the corresponding point within the scene.
Building upon the mentioned concept, we introduce our EvTemMap. Initially, we block all incoming light to the event camera. Then, as we progressively raise the transmittance of the optical system, the brightness of each pixel increases at different rates. Throughout this process, we record the timestamp of IPE for each pixel, forming a Temporal Matrix. Subsequently, the Temporal Matrix undergoes time-intensity mapping through a time-intensity mapping and finally produces a grayscale image. However, due to hardware constraints, the time-intensity mapping is not as ideal as we derive. Various factors such as uncertain contrast threshold, timestamp error, etc. degrade the final image. To address this issue, we design a temporal and spatial degradation space, especially for event-to-image conversion. Trained on the designed degradation model, our temporal-mapping network converts Temporal Matrix to high-quality, high-grayscale-resolution, and high-dynamic-range image faithfully. Additionally, we record Temporal Matrix across diverse scenes and compile a dataset named TemMat dataset. Through extensive comparisons with existing event-to-video methods, we demonstrate the superiority of the proposed EvTemMap.
@@
  • We revisit the mechanism of the event camera, and propose event-based temporal-mapping photography (EvTemMap) to convert events from a stationary event camera in static scenes to image, which is the first work in this area to the best of our knowledge.
  • We further explore the temporal and spatial degradation, design a degradation model, and train a temporal mapping network for our EvTemMap.
@@ 2.1 Event to Video Reconstruction
Non-learning-based methods. The pioneering work [9] demonstrated the feasibility of extracting intensity information from event data and reconstructing images from events captured by an event camera traversing a stationary scene. Their approach mainly depends on the inherent connection between intensity gradients and optical flow, utilizing each event to formulate equations for brightness constancy. Subsequently, event integration was proposed to be independent of assumptions regarding scene structure or motion dynamics [28] by Munda et al.. Scheerlinck et al. [34] defined the reconstruction as an energy minimization problem and applied a high-pass filter to events before integration. However, these methods frequently encounter "bleeding edges" caused by spatially variant contrast thresholds.
Learning-based Methods With the advancement of deep learning, Barua et al. [4] utilized K-SVD [2] on simulated data to train a dictionary mapping small patches of integrated events to image gradients. Subsequently, they employed Poisson integration for intensity image reconstruction. Rebecq et al. [33] designed an end-to-end recurrent model trained with synthetic events, resulting in the high-speed video reconstruction method E2VID. Subsequent works [35, 36, 6, 47] maintained this paradigm while improving efficiency [35], fine-grained details [6], and performance in low-light conditions [47]. More recently, Paredes-Vall’es et al. [30] extended E2VID to self-supervised learning by introducing optical flow as a constraint. However, all these methods only recover intensity up to an unknown initial image, leading to “ghosting” effects and non-uniform grayscale distribution. Moreover, these methods are primarily applicable in dynamic scenes, where either the camera or the scene exhibits motion.
@@ 3.3 Degradation Model for EvTemMap
Cold and hot pixels. Because of the limitation of the circuit, there are cold pixels, which are less prone to generating events, and hot pixels, which constantly generate events in DVS [15]. In the Temporal Matrix, the cold pixel is the location where $t^{*}$ is null, which is artificially set to the maximum value of $t^{*}$, as shown in Figs. 1(a) and 1(b). In contrast, the hot pixel exhibits an unusually early $t^{*}$, resulting in the brightest white dots in the grayscale image, as shown in Fig. 3b. Experimentally, it is found that hot pixels consistently occupy specific locations. We locate the position of hot pixels with pixel calibration [42]. We then remove the hot pixels by defining the $t^{*}$ of them as the minimum of the other positions in Temporal Matrix, i.e. temporal pre-processing.
@@ 4.1 TemMat Dataset
For qualitative comparison, we collect TemMat dataset. The dataset features (1) diverse scenarios including both outdoor and indoor scenes, (2) a significant luminance range from 0.02 lux to 20000 lux, and (3) high dynamic range scenes, reaching up to 114 dB. We collect the Temporal Matrix with AT-DVS. To compare with other E2VID [33, 36] methods, for each scene, we also collect motion events by slowly and gently shaking the DVS, without changing the transmittance of the optical system. For capturing reference images, a Nikon D780 SLR camera equipped with an AF-S NIKKOR 50mm f/1.4G lens, configured in bracket-exposure HDR mode, is employed. To achieve a depth of field comparable to that of the event camera, an aperture of F16 is utilized in aperture priority mode. Because the Temporal Matrix, motion events for E2VID, and HDR images by SLR camera are obtained separately, the results of EvTemMap, E2VID, and reference images are not pixel-aligned.
@@ 4.3 Comparisons with State-of-the-Art Methods
Figure 4 presents qualitative results on the TemMat dataset, comparing our approach with two event-to-video methods: E2VID [32] and E2VID+ [36]. Our approach achieves grayscale image recovery closest to the reference, showcasing the realism in our grayscale restoration in both indoor and outdoor scenes. Specifically, in the leftmost column, our method effectively captures the grayscale transition from outdoors to indoors, and even reflections from glass are discernible. In addition, in the night scenes presented in the two right columns, the trees in the shadows and the lights in the bright areas are correctly grayscale mapped. However, while the E2VID and E2VID+ methods roughly depict the outlines of the scene in the left three columns, their grayscales are significantly distorted. In the last two columns, they recover a grayscale that is even opposed to the real situation.
@@
Figure 5: Quantitative comparison with E2VID using a standard resolution board. The corresponding evaluation score is under each image. EvTemMap achieves the highest line pair evaluation score. Best viewed on a screen and zoomed in to avoid spectral aliasing.
For a quantitative comparison of EvTemMap and the E2VID method regarding texture detail, we use a standard resolution board as the subject and conduct a user study with up to 86 participants. As Fig. 5 shows, participants are asked to identify the highest-grade line pair they can distinguish in each image. The quantitative results under each image demonstrate that our method excels in discerning finer texture details. In dark, medium, and bright environments, our spatial resolution is improved by 17.8%, 191%, and 732% respectively compared to the E2VID method. Note that it is better to zoom in for observation to avoid spectral aliasing.
Compared with the EvTemMap result, the raw intensity image from Temporal Matrix presents more random noise, reducing the contrast of the line pair, resulting in a decrease in line pair resolution.
@@ 4.4 High Dynamic Range Photography
To test the dynamic range of our EvTemMap, we construct an HDR scene with a range of 114dB in a dark room, featuring 20430 lux at the brightest point and 0.04 lux at the darkest point. Fig. 5(a) shows that EvTemMap records details in both shadow and highlight areas without overexposure or underexposure. Around the brightest point, we can observe the gradation of brightness within the elliptical light source as well as the diffracted asterisms around it. Near the darkest point, the result of EvTemMap effectively captures intricate details in the hair of David’s sculpture and natural variations in light and shadow on the shoulder. In the rest of the image, the “ECCV2024” texts on the screen and in the shadows, as well as the background resolution board, are clearly visible. Comparing Fig. 5(a) and Fig. 5(b), it can be found that EvTemMap can truly depict various details in the scene, while E2VID can only restore textures with high contrast. Compared with single-exposure results in Fig. 5(c), our method captures more details in both shadow and highlight grayscale variations than the conventional LDR image. The outcome from the conventional camera using bracketed HDR is depicted in Fig. 5(d) for reference.
@@ 4.6 Motion Photography
(a) EvTemMap
(b) E2VID
(c) Conventional LDR
Figure 8: Comparison in motion scenes. (a) Motion blur in EvTemMap occurs only on one side. (b) Motion blur in E2VID exhibits “ghosting” effects. (c) Motion blur in conventional image occurs on both sides.
Because the time for capturing Temporal Matrix is not ignorable, EvTemMap also experiences motion blur in motion photography, but it presents different characteristics than motion blur in conventional and E2VID images. Fig. 8 shows results from EvTemMap, E2VID, and a traditional camera. Comparing Fig. 7(a) with Fig. 7(c), the motion blur in the EvTemMap image occurs on a single side of the circle during the motion from blue to green, while the motion blur in conventional image occurs on both sides of the circle. Result from E2VID is depicted in Fig. 7(b), which suffers from “ghosting” effects, (c.f. Sec. 2.1). Besides, when event camera remains stationary, E2VID fails to recover the static background.
EvTemMap’s unique motion blur stems from its temporal mapping principle. Different from integral imaging from conventional cameras, where the entire image is captured during a fixed integration process, EvTemMap gets the grayscale by recording the rate of change of brightness of different pixels. Such brightness changes are caused by a combination of transmittance changes and feature motion in motion scenes.
4.7 Downstream Computer Vision task
(a) EvTemMap
(b) E2VID
(c) Conventional LDR
(d) Bracketed HDR
Figure 9: Image segmentation results for each method. (a) presents a satisfactory segmentation result akin to (d). (b): part of the screen is segmented correctly, no target is segmented in the background. (c): The background segmentation fails due to underexposure, but the segmentation results are good in well-lit areas.
Reconstructed images from EvTemMap versatile for various computer vision tasks, which bridges the gap between modern computer vision algorithms and event cameras. Taking image segmentation [48] as an example, we compare the segmentation maps from the images from EvTemMap, E2VID, conventional SLR camera. As shown in Figure 9, EvTemMap, akin to bracketed HDR, achieves outstanding object segmentation in high dynamic range scenes. E2VID fails to segment the detail of the background resolution board due to the large amount of texture missing. The ambiguous grayscale of the floor results in incorrect segmentation. For the conventional LDR image, the underexposed background also fails the segmentation task.
Additional noteworthy results for downstream vision tasks can be explored in the supplementary material.

### v2

@@ Abstract
Event cameras, or Dynamic Vision Sensors (DVS) are novel neuromorphic sensors that capture brightness changes as a continuous stream of “events” rather than traditional intensity frames. Converting sparse events to dense intensity frames faithfully has long been an ill-posed problem. Previous methods have primarily focused on converting events to video in dynamic scenes or with a moving camera. In this paper, for the first time, we realize events to dense intensity image conversion using a stationary event camera in static scenes with a transmittance adjustment device for brightness modulation. Different from traditional methods that mainly rely on event integration, the proposed Event-Based Temporal Mapping Photography (EvTemMap) measures the time of event emitting for each pixel. Then, the resulting Temporal Matrix is converted to an intensity frame with a temporal mapping neural network. At the hardware level, the proposed EvTemMap is implemented by combining a transmittance adjustment device with a DVS, named Adjustable Transmittance Dynamic Vision Sensor (AT-DVS). Additionally, we collected TemMat dataset under various conditions including low-light and high dynamic range scenes. The experimental results showcase the high dynamic range, fine-grained details, and high-grayscale resolution of the proposed EvTemMap. The code and dataset are available in https://github.com/YuHanBaozju/EvTemMap.
@@
Dynamic Vision Sensor Image reconstruction High dynamic range imagingLow-light photography
@@ 1 Introduction
Event cameras offer high temporal resolution and dynamic range, swiftly responding to brightness changes through positive and negative events. While the distinctive asynchronous data output from event cameras diminishes data redundancy, it concurrently discards the grayscale information of the image. To capture grayscale images, a dynamic and active pixel vision sensor (DAVIS) [5] integrates an active pixel sensor (APS) and a DVS sensor on a single chip. This integration allows the sensor to output both grayscale frames and events. The fusion of grayscale images with events facilitates various computer vision tasks. Many event-based image/video enhancement tasks rely on such devices, such as motion deblur [40, 52, 17, 26], video interpolation [41, 43], high dynamic range (HDR) imaging [30], auto-focusing [3] etc. These tasks leverage the high temporal resolution and dynamic range of events to enhance grayscale images.
@@
Since the temporal dimension of events exhibits theoretically infinite description levels, with 1 $\mu$s representing one level, can we utilize time to measure grayscale? Imagine a scene: the event camera is initially in complete darkness, and when the lighting gradually increases, the event camera captures positive events as each pixel gets brighter. Each pixel lights up at a different rate, depending on its corresponding point’s light intensity in the scene. As the event sensor initiates from darkness and gradually intensifies, the time each pixel takes to reach a specified threshold (i.e., the timestamp of the initial positive event, IPE) indicates its brightening speed. Hence, through the measurement of the timestamp associated with each IPE, we can derive the light intensity of the corresponding point within the scene.
Building upon the mentioned basic concept, we introduce our EvTemMap. Initially, we block all incoming light to the event camera. Then, as we progressively raise the transmittance of the optical system, the brightness of each pixel increases at different rates. Throughout this process, we record the timestamp of IPE for each pixel, forming a Temporal Matrix. Subsequently, the Temporal Matrix undergoes time-intensity mapping and finally produces a grayscale image. However, due to hardware constraints, the time-intensity mapping is not as ideal as we derive. Various factors such as uncertain contrast threshold, timestamp error, etc. degrade the final image. To address this issue, we design a temporal and spatial degradation space, especially for event-to-image conversion. Trained on the designed degradation model, our temporal-mapping network converts Temporal Matrix to high-quality, high-grayscale-resolution, and high-dynamic-range image faithfully. Additionally, we record Temporal Matrix across diverse scenes and compile a dataset named TemMat dataset. Through extensive comparisons with existing event-to-video methods, we demonstrate the superiority of the proposed EvTemMap.
@@
  • We revisit the mechanism of the event camera, and propose event-based temporal-mapping photography (EvTemMap)to convert events from a stationary AT-DVS in static scenes to an image, which is the first work in this area to the best of our knowledge.
  • We further explore the temporal and spatial degradation, design a degradation model and train a temporal mapping network for our EvTemMap.
@@ 2.1 Event to Video Reconstruction
Non-learning-based methods. The pioneering work [9] demonstrated the feasibility of extracting intensity information from event data and reconstructing images from events captured by an event camera traversing a stationary scene. Their approach mainly depends on the inherent connection between intensity gradients and optical flow, utilizing each event to formulate equations for brightness constancy. Subsequently, event integration was proposed to be independent of assumptions regarding scene structure or motion dynamics [31] by Munda et al. Scheerlinck et al. [37] defined the reconstruction as an energy minimization problem and applied a high-pass filter to events before integration. However, these methods frequently encounter "bleeding edges" caused by spatially variant contrast thresholds.
Learning-based Methods. With the advancement of deep learning, Barua et al. [4] utilized K-SVD [2] on simulated data to train a dictionary mapping small patches of integrated events to image gradients. Subsequently, they employed Poisson integration for intensity image reconstruction. Rebecq et al. [36] designed an end-to-end recurrent model trained with synthetic events, resulting in the high-speed video reconstruction method E2VID. Subsequent works [38, 39, 6, 51] maintained this paradigm while improving efficiency [38], fine-grained details [6], and performance in low-light conditions [51]. More recently, Paredes-Vall’es et al. [33] extended E2VID to self-supervised learning by introducing optical flow as a constraint. However, all these methods only recover intensity up to an unknown initial image, leading to “ghosting” effects and non-uniform grayscale distribution. Moreover, these methods are primarily applicable in dynamic scenes, where either the camera or the scene exhibits motion.
@@ 3.3 Degradation Model for EvTemMap
Cold and hot pixels. Because of the limitation of the circuit, there are cold pixels, which are less prone to generating events, and hot pixels, which constantly generate events in DVS [15]. In the Temporal Matrix, the cold pixel is the location where $t^{*}$ is null, which is artificially set to the maximum value of $t^{*}$, as shown in Figs. 2(a) and 2(b). In contrast, the hot pixel exhibits an unusually early $t^{*}$, resulting in the brightest white dots in the grayscale image, as shown in Fig. 3b. Experimentally, it is found that hot pixels consistently occupy specific locations. We locate the position of hot pixels with pixel calibration [46]. We then remove the hot pixels by defining the $t^{*}$ of them as the minimum of the other positions in Temporal Matrix, i.e. temporal pre-processing.
@@ 4.1 TemMat Dataset
For comparison, we collect TemMat dataset. The dataset features (1) diverse scenarios including both outdoor and indoor scenes, (2) a significant luminance range from 0.02 lux to 20000 lux, and (3) high dynamic range scenes, reaching up to 114 dB. We collect the Temporal Matrix with AT-DVS. To compare with other E2VID [36, 39] methods, for each scene, we also collect motion events by slowly and gently shaking the DVS, without changing the transmittance of the optical system. For capturing reference images, a Nikon D780 SLR camera equipped with an AF-S NIKKOR 50mm f/1.4G lens, configured in bracket-exposure HDR mode, is employed. To achieve a depth of field comparable to that of the event camera, an aperture of F16 is utilized in aperture priority mode. Because the Temporal Matrix, motion events for E2VID, and HDR images by SLR camera are obtained separately, the results of EvTemMap, E2VID, and reference images are not pixel-aligned.
@@ 4.3 Comparisons with State-of-the-Art Methods
Qualitative results. Qualitative results on the TemMat dataset are presented in Fig. 4, comparing our approach with two event-to-video methods: E2VID [35] and E2VID+ [39]. Our approach achieves grayscale image recovery closest to the reference, showcasing the realism in our grayscale restoration in both indoor and outdoor scenes. Specifically, in the leftmost column, our method effectively captures the grayscale transition from outdoors to indoors, and even reflections from glass are discernible. In addition, in the night scenes presented in the two right columns, the trees in the shadows and the lights in the bright areas are correctly grayscale mapped. However, while the E2VID and E2VID+ methods roughly depict the outlines of the scene in the left three columns, their grayscales are significantly distorted. In the last two columns, they recover a grayscale that is even opposed to the real situation.
@@
Table 1: Quantitative comparison on TemMat dataset.
Methods EvTemMap E2VID [35] E2VID+ [39] SPADE [6]
PSNR 15.75 10.48 11.21 10.20
SSIM 0.69 0.58 0.55 0.52
Quantitative results. Table 1 presents the quantitative comparison results with the E2VID-based methods [35, 39, 6]. We take the bracketed HDR grayscale images captured by the SLR camera as the ground truth, and use LightGlue [27] to align the results of various methods with the ground truth, and then perform PSNR and SSIM evaluations. In terms of quantitative results, EvTemMap performs the best, while the three E2VID-based methods do not differ much. More detailed quantitative and qualitative results can be explored in the supplementary material.
@@ 4.4 High Dynamic Range Photography
To test the dynamic range of our EvTemMap, we construct an HDR scene with a range of 114dB in a dark room, featuring 20430 lux at the brightest point and 0.04 lux at the darkest point measured by an illuminance meter. Figure 5(a) shows that EvTemMap records details in both shadow and highlight areas without overexposure or underexposure. Around the brightest point, we can observe the gradation of brightness within the elliptical light source as well as the diffracted asterisms around it. Near the darkest point, the result of EvTemMap effectively captures intricate details in the hair of David’s sculpture and natural variations in light and shadow on the shoulder. In the rest of the image, the “ECCV2024” texts on the screen and in the shadows, as well as the background resolution board, are clearly visible. Comparing Fig. 5(a) and Fig. 5(b), it can be found that EvTemMap can truly depict various details in the scene, while E2VID can only restore textures with high contrast. Compared with the single-exposure result in Fig. 5(c), our method captures more details in both shadow and highlight grayscale variations than the conventional LDR image. The outcome from the conventional camera using bracketed HDR is depicted in Fig. 5(d) for reference.
@@ 4.6 Motion Photography
Figure 7: Comparison in motion scenes.(a) Motion blur in EvTemMap occurs only on one side. (b) Motion blur in conventional image occurs on both sides. (c) Motion blur in E2VID exhibits “ghosting” effects.
Because the time for capturing Temporal Matrix is not ignorable, EvTemMap also experiences motion blur in motion photography, but it presents different characteristics than motion blur in conventional and E2VID images.
Figure 7 shows results from EvTemMap, a traditional camera and E2VID. The top row shows an example of scene motion, where the fan rotates at a fixed speed. Comparing EvTemMap images with conventional LDR images, the motion blur in the EvTemMap image occurs on a single side of the circle during the motion from blue to green, whereas the motion blur in conventional image occurs on both sides of the circle. In another example of camera motion, AT-DVS and a conventional SLR camera are combined into a coaxial system with a beam splitter to capture a fixed screen while the system moves. With both cameras set to an exposure time of 40 ms, the bottom row demonstrates that EvTemMap produces less motion blur compared to the conventional LDR image. In Fig. 7c, while E2VID also reconstructs grayscale for motion scenes, it suffers from "ghosting" effects (cf. Sec. 2.1) and fails to handle static areas of the scene effectively (e.g., the static background in the top row).
EvTemMap’s unique motion blur characteristics arise from its temporal mapping principle. Unlike conventional cameras that use integral imaging to capture the entire image during a fixed integration period, EvTemMap varies its grayscale acquisition speed based on the brightness of different areas in the image. Brighter areas are captured faster, resulting in less motion blur. Theoretically, EvTemMap’s maximum motion blur occurs in the darkest areas. Although these areas experience the same amount of motion as in conventional images for the same exposure time, the motion blur is only half as prominent because it occurs on only one side, which demonstrates the advantage of EvTemMap over conventional SLR in motion photography.
@@ 5 Conclusion
Acknowledgements
This work was supported in part by the National Key R&D Program of China, under Grant No.2022YFF0705500, in part by the National Natural Science Foundation of China under Grant No.12174341.